# Ambitions for theory in the physics of life

### v1

@@ 1.1 An example, and a problem
Figure 1: Segmented body plans of larval insects, and the underlying molecular blueprint. Top: the caterpillar of the tiger moth, in which the segments are especially visible. Second from top: larva (maggot) of the fruit fly Drosophila melanogaster. Second from bottom: the “cuticle preparation” of the Drosophila maggot shortly after hatching, highlighting the segmented structure [9]. Bottom: An optical section through an embryo stained for three of the “pair rule” proteins, showing striped patterns that align with the body segments; data from Ref [10], with thanks to M Nikolić for the image. Upper panels from R. Tomlinson, Henderson State University, and https://bugguide.net/node/view/1308194/bgimage, respectively.
@@
How does the fly embryo make stripes? By the turn of this century, there was a clear outline of how this works. To start, all the relevant molecules were identified, and this was one of the greatest triumphs of using genetic methods to dissect a complex phenomenon in living systems [9, 11, 12]. The molecules with striped patterns of concentration are a group of eight proteins that are encoded by the “pair rule” genes. Whether these proteins are synthesized [^2] When the information encoded in a particular gene is read out to make the corresponding protein we say that the gene is “expressed.” This occurs in two steps, the synthesis of mRNA from the DNA template (transcription) and then the synthesis of protein from the mRNA template (translation). Both steps are regulated, although our emphasis will be on the regulation of transcription. When someone talks about the “expression level of a gene” it can be ambiguous whether this means the number of mRNA molecules or the number of protein molecules. is controlled largely by the concentrations of another set of four proteins that are encoded in the “gap genes,” so named because a mutation in one of these genes causes a large gap in the body plan. The gap genes not only regulate the pair rule genes, they also regulate one another, forming a network. Finally, there are three inputs to the gap gene network that are provided by the mother when she makes the egg. This flow of information through three layers of a molecular network is schematized in Fig 2.
Figure 2: A schematic of information flow in the early fly embryo. The maternal inputs have simple spatial profiles: one with high concentration at the anterior end (left), one with high concentration at the posterior end (right), and one that is symmetrically high at the both ends (middle). These molecules activate the expression of four gap genes, which also regulate one another. Finally the gap genes modulate the expression of eight pair rule genes, whose concentration profiles consist of striped patterns; one example is shown. Not illustrated are paths by which the maternal inputs can reach around the gap genes to regulate the pair rule genes directly. Images from data in Ref [10], with thanks to MD Petkova.
Figure 3: Parameters hiding under the arrow. $A$ is a transcription factor activating the expression of $B$, as in Eq (1). If this system comes to steady state then the expression level $B=r_{\rm max}\tau f(A)$, and we are free to choose units such that $r_{\rm max}\tau$, so that $B=f(A)$ as shown here. The regulatory function $f(A)$ has at least two parameters, as shown. The system also is noisy, which we neglect in Eq (1), but this will be important below.
@@
  • In the embryo there are many cells, [^4] In the fly, there are actually no membranes separating the cells until well into the fourteenth cycle. It would be more precise to say “many nuclei.” so there is a separate copy of these dynamics at each of $\sim 2^{14}$ sites. Exchange of molecules between sites might be described by an effective diffusion constant, but the dynamics could be more complicated.
@@
Our most complete theories of the natural world certainly have parameters [13], but there is a sense that if we are focused too much on these parameters then we are doing something wrong. If parameters proliferate, we take this as a sign that we are missing some additional level of unification that could relate these many parameters to one another; if our qualitative explanation of phenomena hinges on precise quantitative adjustment of parameters, then we search for the hidden dynamics that could make this apparent fine tuning happen more naturally. Some of the greatest triumphs of modern theoretical physics are nearly free from parameters—the BCS theory of superconductivity [14], the renormalization group theory of critical phenomena [15, 16, 17, 18], the theory of the fractional quantum Hall effect [19], and more. Importantly, these examples refer not to a rarefied world of interactions among small numbers of elementary particles, but rather to the properties of real, macroscopic materials, with all their chemical complexities. In fact the periodic table of the elements, and hence most of the rules of chemical bonding, can be derived from quantum mechanics by knowing only one parameter ($\alpha\sim 1/137$), and even this isn’t so important in the first rows of the table. Perhaps ironically, the standard model of elementary particles has many more parameters [13], but non–trivial predictions that are the foundation for our confidence in this model can be nearly parameter–free, as with the connection of deep inelastic scattering experiments to the asymptotic freedom of QCD [20, 21, 22, 23].
@@ 1.2 Reacting to complexity
A positive way of saying this is that living systems really must be described by models with many parameters, as in the example of the fly embryo, and so the interesting theoretical problems concern how we infer these parameters from data, and how prediction may be possible even when parameters are underdetermined. This once seemed pessimistic, but of course it has received renewed attention in response to the dramatic successes of deep neural networks [24] and large language models [25, 26]. These models grew, over a long period, out of efforts in the physics community to build theories of brain function [27, 28, 29, 30, 31], and many physicists now are interested in the question of why these networks ‘work’ as well as they do [32, 33, 34, 35, 36]. It is not unreasonable to think that a theory of deep networks will circle back to influence how we think about the physics of life.
@@
One sometimes hears the claim that biology is different from physics because biology is historical and physics is not. This is a complicated claim, [^7] In our modern understanding, the particular locations of stars or galaxies is an accident of history, but the distribution out of which these positions are drawn is not. Thus one can see cosmology as historical. But this absolutely does not preclude having theories in the same was as in other areas of physics—based on fundamental principles and tested by detailed comparison to (very!) quantitative experiments. but it suggests that even if we can’t have a theory of life as we see it today, we could have a quantitative theory of the relationships among different life forms, over time—evolution. Indeed, circa 2000, a number of physicists realized that population genetics and evolutionary dynamics could be seen as statistical physics problems, and this has been extraordinarily productive: even the simplest models of evolutionary change have subtle properties, the progress of a population of organisms over a fitness landscape is dominated by individuals in the tail of the distribution [40, 41], and more realistic contexts lead to fascinating interacting many–body problems [42, 43, 44]. This work has involved both sophisticated theory and quantitative connections to experiment, both in the laboratory [45, 46, 47] and in the populations of viruses that infect humans all over the world [48, 49, 50]. [^8] As result there have been important practical consequences of this work, both in designing next season’s flu vaccines and in the global response to covid–19. See, for example, https://nextstrain.org.
@@
Biologists often complain that physicists oversimplify when we think about living systems, but we also simplify when we think about inanimate matter. These simplifications work not just because we are lucky. The renormalization group teaches us that if we start with a detailed microscopic description of a system and coarse–grain to arrive at a model for behavior on long distances and long time scales, then in this process many of the microscopic details will be lost. In technical language, there usually are only a small number of relevant operators [15, 16, 17, 18]. Thus, quantitative descriptions of macroscopic phenomena can be simpler and more universal that the underlying microscopic mechanisms. This inspires us to think that essentially macroscopic functional phenomena in living systems could be similarly independent of molecular details. Possibly related ideas of simplification arise from thinking about the broad spectrum of sensitivities to different parameters [52, 51], and the universal behavior of dynamical systems near transition or bifurcation points [53].
@@ 1.3 A few words about experiments
The relevant variables in a genetic network are the concentrations of proteins and mRNAs. They are ways of measuring both, and this can be done in live embryos and in fixed embryos that give us snapshots of the underlying dynamics. Each method has pros and cons, and methods evolve with time. I am writing this just after hearing about (admittedly preliminary) measurements that were not possible a few months before when I gave these lectures in Les Houches. As a theorist you will need to keep up with what can be done experimentally.
@@
Figure 4 shows an example of these measurements in the embryo, for one of the gap genes. The embryo is intact, but the plane of focus is midway through its depth, so cells are arrayed along the rim of the optical slice. One can resolve individual nuclei, but here we just plot the fluorescence intensity averaged over a window with diameter equal to that of the nuclei, sliding along the rim; we measure the position $x$ of the window by projecting onto the midline, so that $x=0$ is the anterior end (future head) and $x=L$ is the posterior end (future tail). It is hard to convert the raw fluorescence intensity into absolute protein concentrations, so we choose units where the maximum mean concentration across embryos is $\max_{x}\langle g(x/L)\rangle=1$, and we subtract a background so that $\min_{x}\langle g(x/L)\rangle=0$, where $\langle g(x/L)\rangle$ is the average across the ensemble of embryos at fixed $x/L$ [^11] These experiments are done in inbred laboratory stocks of flies, minimizing genetic variation. Nonetheless there are small variations in the embryo length $L$, even in eggs laid by the same mother in succession. Our (very conventional) choice to plot always vs $x/L$ suggests that the underlying pattern formation dynamics have some mechanism that compensates for these variations, achieving a kind of scale–invariance. This is a subtle problem that I stayed away from in my lectures, but we have come back to it since then [58]. To be clear, this single normalization is applied to all the embryos in the sample. Plausible alternatives, such as normalizing each image by the maximum intensity that we see, are unphysical and distort our estimates of noise, which will be important below. The small fluctuations from embryo to embryo that we see here provide a first glimpse of how the picture of biology as noisy can be conquered by careful experiments.
Figure 4: Measuring protein concentrations in the Drosophila embryo. (A) Image of fluorescence from labeled antibodies against one the proteins encoded by the gap genes, corresponding to $g_{2}$ in Fig 11 below; the gene is named krüppel. (B) Raw fluorescence intensity from images of many such embryos, measured along the (straigher) upper edge. Embryos are chosen to be a in narrow time window after the egg is laid, and in a narrow range of orientations. We define a background (lower dashed line) and a scale (upper dashed line) so that the mean concentration is in the range $10<\langle g_{2}(x)\rangle<$ 1. (C) Mean and standard deviation across this large ensemble of embryos. Note that fluctuations in the peak concentration are $\sim 10\%$, and the flanks around the peak are defined very sharply. Data from Ref [10].
@@
All of the proteins we are interested in here are present at very low concentrations, so the proportionality of GFP fluorescence to its concentration is guaranteed. With immunoflourescent staining it is a bit less obvious, because nonlinearities might creep in through the two steps of antibody binding. By making a GFP fusion with a particular protein and then directing immunoflourescent stains against both the protein and GFP itself, one can verify that immunofluorescence intensity really is proportional to protein concentration, and that nothing funny happens to break the 1–to–1 link between the protein and the GFP tag [64].
The construction of GFP fusion proteins has the obvious advantage that one measure protein concentrations in live cells or embryos, making the dynamics of these signals (literally) visible in real time, The disadvantage is that GFP does not immediately fold into the structure that supports maximal fluorescence, but rather takes time to “mature.” There is a continuing stream of work to engineer GFP variants that have shorter maturation times, but we are not quite where we would like to be. Thus, if we want to measure the concentration of a maternal morphogen during nuclear cycle fourteen, we are looking at proteins that have been synthesized over more than two hours and everything is fine. [^12] Dual immunofluorescence experiments (above) can detect small corrections due to the maturation time. On the other hand, if we make GFP fusions with the pair rule genes we can see the stripes emerging during cycle fourteen but the dynamics we see probably are lagging the true dynamics.
@@
Figure 5: Single molecule mRNA detection in a small region of the Drosophila embryo. Green probes are complementary to the mRNA of the hb gene (protein concentration is shown as $g_{1}$ in Fig 11) and all nuclei are labeled in magenta. Detailed analysis shows that all the spots outside the nuclei have brightness drawn from a narrow distribution, and a size consistent with the point spread function of the microscope; these and other observations indicate that they are single molecules. The one or two brighter spots inside the nuclei are point at which mRNAs are being transcribed, as explained in the text. Expression is larger toward the left, corresponding the anterior of the embryo. Scale bar $5\,\mu{\rm m}$. From Ref [66], with thanks to T Gregor.
In Figure 5 you see not only the individual mRNA molecules in the cytoplasm, but also one or two exceedingly bright spots inside the nuclei. These are the locations along the two chromosomes where the mRNA for this gene is being transcribed. The gene is long enough that many copies of the transcriptional apparatus can operate simultaneously, with the result that many mRNA strands are “in progress” and still tethered to the DNA. With care one can calibrate against the fluorescence intensity of the cytoplasmic spots and effectively count these nascent transcripts. By tagging probes directed against the early and late parts of the sequence one can gets a distribution of both colors and intensities across spots, and this can be used to test models of the underlying dynamics [69]. By interleaving the different colors one can check the precision of the measurement.
As with proteins, there is a strategy for live measurements of transcriptional activity. Instead of attaching the DNA sequence encoding GFP to the gene of interest, one can add a sequence drawn from a virus that carries its genome as RNA rather than DNA. To package the genome, proteins that form the coat of the virus bind to these specific RNA sequences, which fold into three–dimensional structures called “stem loops.” If the fly also has been genetically engineered to produce the coat protein tagged with GFP, then as the gene we are interested in gets transcribed these molecules, initially distributed throughout the nucleus will bind to the nascent transcript [70, 71, 72]. If there are many stem loops there will be many GFPs, enough to light up the mRNAs much as in the nuclear spots of Fig 5. After a decade of development the noise levels in these measurements now are at the point where one can almost count transcripts one by one [73].
@@
Figure 6: Recording action potentials from retinal ganglion cells. (A) Salamander retina on an array of electrodes. The electrodes, and the leads that carry signals away from the electrodes, are black features on a transparent slide; the ganglion cells of the retina have been filled with a green dye. Round green objects are cell bodies, and long lines are bundles of axons that eventually converge to form the optic nerve. Note that the number of electrodes is comparable to the number of cells (B) The next generation of electrode arrays. (C) Voltage traces from a selection of these electrodes during an experiment on the salamander retina. Blue traces are the actual voltages, and orange traces are a reconstruction of the voltages as a superposition of stereotyped waveforms—action potentials (spikes) from individual neurons—learned from a different part of the data. (D) Raster plot of the action potentials found in (C) for $100+$ cells, a dot showing the time of a spike from each neuron. (E) Expanded version of the red box, with responses expressed as binary (spike/silence) words in $20\,{\rm ms}$ bins. From experiments by MJ Berry II and colleagues, including D Amodei, O Marre, JL Puchalla, and R Segev, with thanks [80, 82].
@@
Figure 7: Monitoring electrical activity in the brain by imaging of calcium sensitive fluorescent proteins as a mouse moves in virtual reality. (A) Schematic of the experiment, with mouse (in black) running on a styrofoam ball. Motion of the ball advances the position of a virtual world projected on a surrounding toroidal screen, and a scanning two–photon microscope is focused on a single layer of cells in the hippocampus. (B) Fluorescence image of these neurons expressing calcium sensitive fluorescent protein. (C) Integrated fluorescence from a single cell vs time. Note the very quiet baseline interrupted by discrete events. These can be binarized, and (as shown in the insets) we understand enough about the dynamics of the indicator molecule to assign unusually slow transient decays as flickering off and on. From experiments by JL Gauthier, CD Brody, and DW Tank, with thanks [105, 106].
@@
For smaller animals, such as the worm C elegans or the larval zebrafish, we are getting close to recording from every single neuron in the brain at reasonable time resolution [99, 100]. The same techniques of genetic engineering that have led to calcium–sensitive fluorescent proteins have also led to proteins that act as light–sensitive ion channels [101, 102, 103]. Combining these tools means that one can both inject currents into neurons and records their responses, all using light and working at single cell resolution. In C elegans this has led to a nearly complete “pump–probe” experiment probing signal transfer among all $\sim 10^{4}$ pairs of neurons [104].

### v2

@@ 1.1 An example, and a problem
Figure 1: Segmented body plans of larval insects, and the underlying molecular blueprint. (A) Caterpillar of the agreeable tiger moth, in which the segments are especially visible. Image by Cyndy Syms Parr, from Wikipedia under the CC–SA 2.0 license. (B) Larva (maggot) of the fruit fly Drosophila melanogaster. Image by Salvatore Vitanza, with permission. (C) The “cuticle preparation” of the Drosophila maggot shortly after hatching, highlighting the segmented structure. Thanks to Eric Wieschaus for the image, from experiments described in Ref [9]. (D) An optical section through an embryo stained for three of the “pair–rule” proteins, showing striped patterns that align with the body segments; data from Ref [10], with thanks to M Nikolić for the image.
@@
How does the fly embryo make stripes? By the turn of this century, there was a clear outline of how this works. To start, all the relevant molecules were identified, and this was one of the greatest triumphs of using genetic methods to dissect a complex phenomenon in living systems [9, 11, 12]. The molecules with striped patterns of concentration are a group of eight proteins that are encoded by the “pair–rule” genes. Whether these proteins are synthesized [^2] When the information encoded in a particular gene is read out to make the corresponding protein we say that the gene is “expressed.” This occurs in two steps, the synthesis of mRNA from the DNA template (transcription) and then the synthesis of protein from the mRNA template (translation). Both steps are regulated, although our emphasis will be on the regulation of transcription. When someone talks about the “expression level of a gene” it can be ambiguous whether this means the number of mRNA molecules or the number of protein molecules. is controlled largely by the concentrations of another set of four proteins that are encoded in the “gap genes,” so named because a mutation in one of these genes causes a large gap in the body plan. The gap genes not only regulate the pair–rule genes, they also regulate one another, forming a network. Finally, there are three inputs to the gap gene network that are provided by the mother when she makes the egg. This flow of information through three layers of a molecular network is schematized in Fig 2.
Figure 2: A schematic of information flow in the early fly embryo. The maternal inputs have simple spatial profiles: one with high concentration at the anterior end (left), one with high concentration at the posterior end (right), and one that is symmetrically high at the both ends (middle). These molecules activate the expression of four gap genes, which also regulate one another. Finally the gap genes modulate the expression of eight pair–rule genes, whose concentration profiles consist of striped patterns; one example is shown. Not illustrated are paths by which the maternal inputs can reach around the gap genes to regulate the pair–rule genes directly. Images reproduced from Ref [10], with permission.
Figure 3: Parameters hiding under the arrow. $A$ is a transcription factor activating the expression of $B$, as in Eq (1). If this system comes to steady state then the expression level $B=r_{\rm max}\tau f(A)$, and we are free to choose units such that $r_{\rm max}\tau=1$, so that $B=f(A)$ as shown here. The regulatory function $f(A)$ has at least two parameters, as shown. The system also is noisy, which we neglect in Eq (1), but this will be important below.
@@
  • In the embryo there are many cells, [^4] In the fly, there are actually no membranes separating the cells until well into the fourteenth cycle. It would be more precise to say “many nuclei.” so there is a separate copy of these dynamics at each of $\sim 2^{14}$ sites. Exchange of molecules between sites can be described by an effective diffusion constant, but the dynamics could be more complicated.
@@
In the traditional core of theoretical physics our models of the world have parameters [13], but somehow fitting these parameters does not seem to be the central focus. We assume that if there are too many parameters there must be something that unifies them, and if we need to set these parameters precisely to non–generic values we hope there is some extra dynamics that can make this happen more naturally. More strongly, their are spectacular successes which are almost parameter independent, such as the BCS theory of superconductivity [14], the renormalization group theory of critical phenomena [15, 16, 17, 18], the theory of the fractional quantum Hall effect [19], and more. These theories make detailed quantitative predictions about the properties of real materials, despite the complexity of these materials on an atomic scale. In fact the periodic table of the elements, and hence most of the rules of chemical bonding, can be derived from quantum mechanics by knowing only one parameter ($\alpha\sim 1/137$), and even this isn’t so important in the first rows of the table. Perhaps ironically, the standard model of elementary particles has many more parameters [13], but non–trivial predictions that are the foundation for our confidence in this model are nearly parameter–free, as with the connection of deep inelastic scattering experiments to the asymptotic freedom of QCD [20, 21, 22, 23].
@@ 1.2 Reacting to complexity
A positive way of saying this is that living systems really must be described by models with many parameters, as in the example of the fly embryo, and so the interesting theoretical problems concern how we infer these parameters from data, or how prediction may be possible even when parameters are underdetermined. This once seemed pessimistic, but it has received renewed attention in response to the dramatic successes of deep neural networks [24] and large language models [25, 26]. These models grew, over a long period, out of efforts in the physics community to build theories of brain function [27, 28, 29, 30, 31], and many physicists now are interested in the question of why these networks ‘work’ as well as they do [32, 33, 34, 35, 36]. It is not unreasonable to think that a theory of deep networks will circle back to influence how we think about the physics of life.
@@
One sometimes hears the claim that biology is different from physics because biology is historical and physics is not. This is a complicated claim, [^7] In our modern understanding, the particular locations of stars or galaxies are accidents of history, but the distribution out of which these positions are drawn is not. Thus one can see cosmology as historical. But this absolutely does not preclude having theories in the same way as in other areas of physics—based on fundamental principles and tested by detailed comparison to (very!) quantitative experiments. but it suggests that even if we can’t have a theory of life as we see it today, we could have a quantitative theory of the relationships among different life forms, over time—evolution. Indeed, circa 2000, a number of physicists realized that population genetics and evolutionary dynamics could be seen as statistical physics problems, and this has been extraordinarily productive: even the simplest models of evolutionary change have subtle properties, the progress of a population of organisms over a fitness landscape is dominated by individuals in the tail of the distribution [40, 41], and more realistic contexts lead to fascinating interacting many–body problems [42, 43, 44]. This work has involved both sophisticated theory and quantitative connections to experiment, both in the laboratory [45, 46, 47] and in the populations of viruses that infect humans all over the world [48, 49, 50]. [^8] As a result there have been important practical consequences of this work, both in designing next season’s flu vaccines and in the global response to covid–19. See, for example, https://nextstrain.org.
@@
Biologists often complain that physicists oversimplify when we think about living systems, but we also simplify when we think about inanimate matter. These simplifications work not just because we are lucky. The renormalization group teaches us that if we start with a detailed microscopic description of a system and coarse–grain to arrive at a model for behavior on long distances and long time scales, then in this process many of the microscopic details will be lost. In technical language, there usually are only a small number of relevant operators [15, 16, 17, 18]. Thus, quantitative descriptions of macroscopic phenomena can be simpler and more universal that the underlying microscopic mechanisms. This inspires us to think that essentially macroscopic functional phenomena in living systems could be similarly independent of microscopic details. Possibly related ideas of simplification arise from thinking about the broad spectrum of sensitivities to different parameters [52, 51], and the universal behavior of dynamical systems near transition or bifurcation points [53].
@@ 1.3 A few words about experiments
The relevant variables in a genetic network are the concentrations of proteins and mRNAs. There are ways of measuring both, and this can be done in live embryos and in fixed embryos that give us snapshots of the underlying dynamics. Each method has pros and cons, and methods evolve with time. I am writing this just after hearing about (admittedly preliminary) measurements that were not possible a few months before when I gave these lectures in Les Houches. As a theorist you will need to keep up with what can be done experimentally.
@@
Figure 4 shows an example of these measurements in the embryo, for one of the gap genes. The embryo is intact, but the plane of focus is midway through its depth, so cells are arrayed along the rim of the optical slice. One can resolve individual nuclei, but here we just plot the fluorescence intensity averaged over a window with diameter equal to that of the nuclei, sliding along the rim; we measure the position $x$ of the window by projecting onto the midline, so that $x=0$ is the anterior end (future head) and $x=L$ is the posterior end (future tail). It is hard to convert the raw fluorescence intensity into absolute protein concentrations, so we choose units where the maximum mean concentration across embryos is $\max_{x}\langle g(x/L)\rangle=1$, and we subtract a background so that $\min_{x}\langle g(x/L)\rangle=0$, where $\langle g(x/L)\rangle$ is the average across the ensemble of embryos at fixed $x/L$ [^11] These experiments are done in inbred laboratory stocks of flies, minimizing genetic variation. Nonetheless there are small variations in the embryo length $L$, even in eggs laid by the same mother in succession. Our (very conventional) choice to plot always vs $x/L$ suggests that the underlying pattern formation dynamics have some mechanism that compensates for these variations, achieving a kind of scale–invariance. This is a subtle problem that I stayed away from in my lectures, but we have come back to it since then [58]. To be clear, this single normalization is applied to all the embryos in the sample. Plausible alternatives, such as normalizing each image by the maximum intensity in that image, are unphysical and distort our estimates of noise, which will be important below. The small fluctuations from embryo to embryo that we see here provide a first glimpse of how the picture of biology as noisy can be conquered by careful experiments.
Figure 4: Measuring protein concentrations in the Drosophila embryo. (A) Image of fluorescence from labeled antibodies against one of the proteins encoded by the gap genes, corresponding to $g_{2}$ in Fig 11 below; the gene is named krüppel. Embryo is $\sim 500\,\mu{\rm m}$ long. Reproduced from Ref [10], with permission. (B) Raw fluorescence intensity from images of many such embryos, measured along the (straigher) upper edge. Embryos are chosen to be a in narrow time window after the egg is laid, and in a narrow range of orientations. We define a background (lower dashed line) and a scale (upper dashed line) so that the mean concentration is in the range $10<\langle g_{2}(x)\rangle<$ 1. (C) Mean and standard deviation across this large ensemble of embryos. Note that fluctuations in the peak concentration are $\sim 10\%$, and the flanks around the peak are defined very sharply. Data from Ref [10], with thanks to T Gregor, MD Petkova, G Tkačik, and EF Wieschaus.
@@
All of the proteins we are interested in here are present at very low concentrations, so the proportionality of GFP fluorescence to its concentration is guaranteed. With immunofluorescent staining it is a bit less obvious, because nonlinearities might creep in through the two steps of antibody binding. By making a GFP fusion with a particular protein and then directing immunofluorescent stains against both the protein and GFP itself, one can verify that immunofluorescence intensity really is proportional to protein concentration, and that nothing funny happens to break the 1–to–1 link between the protein and the GFP tag [64].
The construction of GFP fusion proteins has the obvious advantage that one can measure protein concentrations in live cells or embryos, making the dynamics of these signals (literally) visible in real time. The disadvantage is that GFP does not immediately fold into the structure that supports maximal fluorescence, but rather takes time to “mature.” There is a continuing stream of work to engineer GFP variants that have shorter maturation times, but we are not quite where we would like to be. Thus, if we want to measure the concentration of a maternal morphogen during nuclear cycle fourteen, we are looking at proteins that have been synthesized more than two hours ago and everything is fine. [^12] Dual immunofluorescence experiments (above) can detect small corrections due to the maturation time. On the other hand, if we make GFP fusions with the pair–rule genes we can see the stripes emerging during cycle fourteen but the dynamics we see probably are lagging the true dynamics.
@@
Figure 5: Single molecule mRNA detection in a small region of the Drosophila embryo. Yellow probes are complementary to the mRNA of the hb gene (protein concentration is shown as $g_{1}$ in Fig 11) and all nuclei are labeled in blue. Detailed analysis shows that all the spots outside the nuclei have brightness drawn from a narrow distribution, and a size consistent with the point spread function of the microscope; these and other observations indicate that they are single molecules. The one or two brighter spots inside the nuclei are points at which mRNAs are being transcribed, as explained in the text. Expression is larger toward the left, corresponding to the anterior of the embryo. Scale bar $5\,\mu{\rm m}$. Image courtesy of T Gregor, from the experiments of Ref [66].
In Figure 5 you see not only the individual mRNA molecules in the cytoplasm, but also one or two exceedingly bright spots inside the nuclei. These are the locations along the two chromosomes where the mRNA for this gene is being transcribed. The gene is long enough that many copies of the transcriptional apparatus can operate simultaneously, with the result that many mRNA strands are “in progress” and still tethered to the DNA. With care one can calibrate against the fluorescence intensity of the cytoplasmic spots and effectively count these nascent transcripts. By tagging probes directed against the early and late parts of the sequence one gets a distribution of both colors and intensities across spots, and this can be used to test models of the underlying dynamics [69]. By interleaving the different colors one can check the precision of the measurement.
As with proteins, there is a strategy for live measurements of transcriptional activity. Instead of attaching the DNA sequence encoding GFP to the gene of interest, one can add a sequence drawn from a virus that carries its genome as RNA rather than DNA. To package the genome, proteins that form the coat of the virus bind to these specific RNA sequences, which fold into three–dimensional structures called “stem loops.” If the fly also has been genetically engineered to produce the coat protein tagged with GFP, then as the gene we are interested in gets transcribed these molecules, initially distributed throughout the nucleus, will bind to the nascent transcript [70, 71, 72]. If there are many stem loops there will be many GFPs, enough to light up the mRNAs much as in the nuclear spots of Fig 5. After a decade of development the noise levels in these measurements now are at the point where one can almost count transcripts one by one [73].
@@
Figure 6: Recording action potentials from retinal ganglion cells. (A) Salamander retina on an array of electrodes. The electrodes, and the leads that carry signals away from the electrodes, are black features on a transparent slide; the ganglion cells of the retina have been filled with a green dye. Round green objects are cell bodies, and long lines are bundles of axons that eventually converge to form the optic nerve. Note that the number of electrodes is comparable to the number of cells (B) The next generation of electrode arrays. (C) Voltage traces from a selection of these electrodes during an experiment on the salamander retina. Blue traces are the actual voltages, and orange traces are a reconstruction of the voltages as a superposition of stereotyped waveforms—action potentials (spikes) from individual neurons—learned from a different part of the data. (D) At left, a raster plot of the action potentials found in (C) for $100+$ cells, a dot showing the time of a spike from each neuron. At right, expanded version of the red box, with responses expressed as binary (spike/silence) words in $20\,{\rm ms}$ bins. (A–C) reproduced from Ref [57], with permission; data from experiments by MJ Berry II and colleagues, including D Amodei, O Marre, JL Puchalla, and R Segev, with thanks [80, 82].
@@
Figure 7: Monitoring electrical activity in the brain by imaging of calcium sensitive fluorescent proteins as a mouse moves in virtual reality. (A) Schematic of the experiment, with mouse (in black) running on a styrofoam ball. Motion of the ball advances the position of a virtual world projected on a surrounding toroidal screen, and a scanning two–photon microscope is focused on a single layer of cells in the hippocampus. (B) Fluorescence image of these neurons expressing calcium sensitive fluorescent protein. (C) Integrated fluorescence from a single cell vs time. Note the very quiet baseline interrupted by discrete events. These can be binarized, and (as shown in the insets) we understand enough about the dynamics of the indicator molecule to assign unusually slow transient decays as flickering off and on. (A) and (B) reproduced from Ref [105]; (C) reproduced from Ref [106], with permission. With thanks to L Mehsulam, JL Gauthier, CD Brody, and DW Tank.
@@
For smaller animals, such as the worm C elegans or the larval zebrafish, we are getting close to recording from every single neuron in the brain at reasonable time resolution [99, 100]. The same techniques of genetic engineering that have led to calcium–sensitive fluorescent proteins have also led to proteins that act as light–sensitive ion channels [101, 102, 103]. Combining these tools means that one can both inject currents into neurons and record their responses, all using light and working at single cell resolution. In C elegans this has led to a nearly complete “pump–probe” experiment probing signal transfer among all $\sim 10^{4}$ pairs of neurons [104].